OpenStudy (anonymous):

If 7p + 5q = 2 and 8p - 9q = 17, then -p + 14q =

OpenStudy (johnweldon1993):

*Hint...try subtracting the 2 equations \[7p + 5q = 2\] - \[8p - 9q = 17\] Ill start you off...when you subtract 7p = 8p you get -p *looks like the start of your answer... when you subtract 5q - (-9q) you get 5q + 9q = 14q *looks even better... so to figure out what -p + 14q equals...just subtract 2 - 17.....= ...?

OpenStudy (anonymous):

The answer choices are a. 15 b.19 c.-15 d.-19 e.1

OpenStudy (johnweldon1993):

Right...just follow what I said above... I showed you how to arrive at the equation they wanted....and I've given you the way to calculate the answer 2 - 17 = ...?